OpenStudy (anonymous):

Please help! I give out medals, and I'll be a fan! 1. 10 • [2 + (24 – 8)] ÷ 4 2. Two opposite sides of a square are increased by 5 cm to form a rectangle. If the rectangle has a perimeter of 70 cm, what was the length of a side of the original square?

OpenStudy (johnweldon1993):

1) order of operations...do the innermost parenthesis...(24 - 8) then finish the parenthesis [2 + (what you get for an answer)] after that you go from left to right. what do you get?

OpenStudy (anonymous):

order of operations

OpenStudy (anonymous):

45?

OpenStudy (johnweldon1993):

that is correct if my mental math is good lol

OpenStudy (anonymous):

1) 180/4 = 45 2)

OpenStudy (johnweldon1993):

Okay now 2) Two opposite sides of a square are increased by 5 cm to form a rectangle. If the rectangle has a perimeter of 70 cm, what was the length of a side of the original square? We know a square had 4 equal sides...and a rectangle has 2 equal sides and 2 equal opposite sides... if you increase 2 sides by 5...and now have a perimeter of 70...that is x + x + (x+5) + (x+5) = 70 so that is the same as 4x + 10 = 70 4x = 60 x = ?

OpenStudy (anonymous):

2)60

OpenStudy (johnweldon1993):

2) is not 60...

OpenStudy (anonymous):

15?

OpenStudy (anonymous):

try again?

OpenStudy (anonymous):

@johnweldon1993

OpenStudy (johnweldon1993):

yes 15....it asks what the length of each side WAS....if each side was 15 that would make the perimeter 60...if you add 5 to 2 sides that increases the total perimeter to 70 ...which makes sense! so yes 15 would be your answer

OpenStudy (anonymous):

Ok, thank you so much!!

OpenStudy (johnweldon1993):

does that make sense?

OpenStudy (anonymous):

YES!!!!!!